a circle has diameter 12m.what will be the cost of fencing the circle if the cost is rs.70/m
Answer by fcabanski(1391)

The fence goes around the circle's perimeter. P = pi*d = pi*12 = approx. 37.699m.


The total cost is the price per meter times meters of fencing = 70 * 37.6991 = approx. rs. 2638.94